OpenStudy (anonymous):

Can someone please help me find the foci, vertices, and asymptotes of \[\frac{ (x-4)^2 }{ 36 }-\frac{ (y-2)^2 }{ 9 }=1\]

OpenStudy (anonymous):

@Hero @agent0smith @dan815 can any of you help me with this? :)

OpenStudy (anonymous):

center you read off right from the equation \[\frac{ (x-4)^2 }{ 36 }-\frac{ (y-2)^2 }{ 9 }=1\]looks like \[\frac{ (x-h)^2 }{ a^2 }-\frac{ (y-k)^2 }{ b^2 }=1\]where the center is \((h,k)\) so your center is \((4,2)\) also your \(a^2=36\) and so \(a=6\) also \(b^2=9\) making \(b=3\)

OpenStudy (anonymous):

I have that the center is (4,2) Eccentricity is √3/2 Length of Transverse Axis is 12 Length of Conjugate Axis is 6 but I need help finding the foci, vertices, and asymptotes

OpenStudy (anonymous):

I have no idea what the formulas are for those 3 or how to find them :(

OpenStudy (anonymous):

before we can find the foci, we need to know what it looks like do you know how it is oriented?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):



OpenStudy (anonymous):

like the first one

OpenStudy (anonymous):

ok good that tells you that the vertices are to the right and left of the center (not above and below)

OpenStudy (anonymous):

oh okay, now what do i do? :)

OpenStudy (anonymous):

and now it is easy, because you have \(a^2=36\) making \(a=6\) go \(6\) units to the right and left of \((4,2)\)

OpenStudy (anonymous):

sooo would that be (-6,2) and (6,2)??

OpenStudy (anonymous):

no, 6 units to the right and left of \((4,2)\) not to the right and left of \((0,2)\)

OpenStudy (anonymous):

in simpler english, add and subtract \(6\) to the first coordinate \(4\)

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

oh so (-10,2) and (10,2)?

OpenStudy (anonymous):

not quite \(4+6=10\) so the \((10,2)\) is right

OpenStudy (anonymous):

how about the second one?

OpenStudy (anonymous):

ahh, so would it be -2,2?

OpenStudy (anonymous):

yes it would!

OpenStudy (anonymous):

thanks! so that is the foci?

OpenStudy (anonymous):

now we have the vertices, namely \((-2,2)\) and \((10,2)\) next are the foci

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

no that was the vertices we do the foci now

OpenStudy (anonymous):

again we are going to go the same number of units to the left and right of the center, but this time we need \(c\) which is found via \(c^2=a^2+b^2\) just like in pythagoras since \(a^2=36\) and \(b^2=9\) we have \(c^2=36+9=45\) and so \(c=\sqrt{45}\)

OpenStudy (anonymous):

your teacher probably wants you to write \(\sqrt{45}\) in simplest radical form as \(3\sqrt{5}\)

OpenStudy (anonymous):

yes! like that, so that would be the vertices or is there more to it?

OpenStudy (anonymous):

you have to add and subtract that from the first coordinate of the center, just like we did for the vetices but in this case you just and and subtract by putting a plus and minus sign there

OpenStudy (anonymous):

in other words they look like \[(4-3\sqrt5,2),(4+3\sqrt5,2)\]

OpenStudy (anonymous):

would that be final or do i actually subtract and add that?

OpenStudy (anonymous):

just leave it like that you cannot actually add or subtract unless you take a calculator and find a decimal approximation for those number i would leave them be

OpenStudy (anonymous):

i guess all that is left is the asymptotes right?

OpenStudy (anonymous):

thank you so much! & yes please :)

OpenStudy (anonymous):

for that you need the slope(s)

OpenStudy (anonymous):

and that is also pretty easy, since the slopes are \(\pm\frac{b}{a}\) which in your case is \(\pm\frac{6}{3}=\pm 2\)

OpenStudy (anonymous):

whoa whoa typo there sorry

OpenStudy (anonymous):

it is \(\pm\frac{b}{a}\) all right, but your \(a=6\) and \(b=3\) i had it backwards so your slopes are \(\pm\frac{3}{6}=\pm\frac{1}{2}\)

OpenStudy (anonymous):

its okay! lol

OpenStudy (anonymous):

then the lines must go through the center, \((4,2)\) so by the point slope formula, if the slope is \(\frac{1}{2}\) then the line is \[y-2=\frac{1}{2}(x-4)\] or solving for \(y\) you get \[y=\frac{1}{2}x\]

OpenStudy (anonymous):

if the slope is \(-\frac{1}{2}\) you get \[y-2=-\frac{1}{2}(x-4)\] solve for \(y\) and get \[y=-\frac{1}{2}x+4\]

OpenStudy (anonymous):

if you need another worked out example with a good explanation, google "hyperbola purplemath"

OpenStudy (anonymous):

http://www.purplemath.com/modules/hyperbola2.htm

OpenStudy (anonymous):

so it would just by a vertical asymptote of y=-1/2x+4??

OpenStudy (anonymous):

they do not have vertical asymptotes, they have two slant asymptotes you have to write two of them one is \(y=\frac{1}{2}x\) the other is \(y=-\frac{1}{2}x+4\) both are asymptotes, neither are vertical